During each cycle of its operation, an air conditioner transfers 930 kJ of heat from a room and outputs 1730 kJ of heat to the outdoors. What mechanical work is involved per cycle?

2660 kJ of work performed by the air conditioner

800 kJ of work performed on the air conditioner

2660 kJ of work performed on the air conditioner

800 kJ of work performed by the air conditioner
1
Expert's answer
2018-04-26T10:06:08-0400
W=Q_H-Q_C=1730-930=800 kJ